# Stochastic surfing turbulent vorticity

### v1

@@ Abstract
In the strongly subcritical regime $\omega_{H}/\omega_{a}\ll 1/2$, the resonance peak is purely set by the turbulent noise, making it sensitive to the preferential sampling effect [7, 10, 11] (see Supplementary Information for neutral and heavy particles results). Here we introduce the relaxation time $\tau^{*}=1/\left(\omega_{a}\sqrt{1-4\big{(}\frac{\omega_{H}}{\omega_{a}}\big{)}^{2}}\right)$. Physically, $\tau^{*}$ represents the characteristic time it takes for a perturbed particle rotational motion to return to its phase-locked state, i.e., $\dot{\beta}=0$, after a small disturbance (detailed derivations are in the Supplementary Information). Close to the critical ratio $\omega_{H}/\omega_{a}=1/2$, the significantly increasing particle relaxation time $\tau^{*}$ shifts the resonance position, making it the dominant factor. Therefore, a longer $\tau^{*}$ allows particles to escape the potential well even at lower noise levels.
@@
A striking consequence of SR is the symmetry-breaking effect, manifesting as a “zero plus zero is greater than zero” phenomenon. This leverages the fact that in the weak noise regime $\omega_{\eta}<\omega_{a}$, the response of the system to the noise is non-linear (Fig. 4b, purple-shaded part).
To reveal this effect, instead of applying a constant-frequency rotating magnetic field ($\omega_{H}$), we now impose a time-dependent field rotation with a zero-average, $\omega_{H}(t)$ (Fig. 5a). Each oscillation cycle consists of a forward and a reverse rotation phase, characterized by angular velocities $\omega_{H}^{+}$ and $\omega_{H}^{-}$ with corresponding durations $T_{H}^{+}$ and $T_{H}^{-}$. These parameters are scaled as $\omega_{H}^{+}/\omega_{H}^{-}=T_{H}^{-}/T_{H}^{+}=\gamma$, ensuring that the net angular velocity over a full cycle remains zero. Two distinct response regimes emerge as $\omega_{\eta}/\omega_{a}$ increases for a fixed $\gamma$ (Fig. 5b), highlighting the ability of this oscillation framework to probe the nonlinearity of the system in response to noise. Specifically, the response $\langle\dot{\beta}\rangle/\omega_{\eta}$ decreases monotonically in the nonlinear regime (weak noise limit), while it remains nearly constant and close to zero in the linear regime (strong noise limit, characterized by response collapse, see Fig. 4b, inset). This behavior aligns well with a superposition model which predicts the response as
@@
Conclusion and outlook
We have revealed the first evidence of stochastic resonance (SR) for magnetic particles suspended in turbulence, demonstrating how the interplay between turbulent fluctuations and external forcing induces coherent rotational motion of magnetic particles. Three distinct regimes: “phase-locked”, “back-and-forth”, and “turbulent-dynamics” are identified, with stochastic resonance occurring at the transition between “phase-locked” and “turbulent-dynamics” regimes. Remarkably, even when both turbulence and the magnetic field have zero mean angular velocity, their collective effect is capable of generating a nonzero particle rotational response, as the turbulent vorticity helps the particle to “surf” the rotating magnetic field.
Different particle properties influence their response, and when combined with the stochastic resonance mechanism, they enable a novel magnetic resonance-based approach to measure turbulent vorticity. This method leverages activated magnetic particles as probes, effectively functioning as a turbulent vorticity “microscope” that operates in various conditions. When the magnetic field emitted by the magnetic spinners can be measured directly, it can even be used in optically inaccessible conditions. The protocol is straightforward: by dispersing magnetic particles into turbulence and varying the magnetic frequency $\omega_{H}$ and intensity $\omega_{a}$ while maintaining a constant ratio $\omega_{H}/\omega_{a}$, one can measure the resonance curve of the averaged particle angular response. The location of the resonance peak directly reveals the turbulent vorticity magnitude.
Our findings also open possibilities for flexible manipulation of turbulence, where propeller-like particles could be tuned to navigate specific flow structures, inject energy into turbulence, and influence flow properties. More broadly, our study provides fundamental mechanisms that externally tunable rotational dynamics could be leveraged to design active materials [45, 46, 47]. If organized coherently, their collective behavior may also exhibit hydrodynamic properties analogous to systems with odd viscosity, which may be promising experimental realizations of chiral fluids [48, 49].
@@ Methods
The experimental setup comprises three fundamental components: a turbulence generator, a magnetic field generator, and magnetic particles. Detailed specifications of the experimental apparatus, along with a comprehensive description of the particle rotation tracking methodology, are provided in our dedicated instrumentation and methodology paper. Here, we present only the essential information pertinent to the current study.
@@
The working fluid is water, contained within an octagonal-shaped vessel with an internal diameter of $1502R=$ 150 mm and a height of 220 mm. This specific geometry is chosen to enhance optical accessibility for visualization. The flow is driven by two counter-rotating disks, each with a diameter of 150 mm, separated by a distance of 150 mm, and rotating at the same angular velocity magnitude in opposite directions. Each disk is equipped with six straight blades, each 10 mm in height, designed to enhance inertial stirring. The disks are powered by two calibrated DC motors operating at a constant voltage, ensuring that the angular velocity $\Omega$ remains stable over time with a precision of approximately 1%. The motor rotation frequency can be adjusted within the range of $0.25$ to $8$ Hz. Under these conditions, a fully developed turbulent flow is generated in the central region of the container [38, 39, 50, 40, 51, 52, 41]. This region, measuring approximately $20$ mm $\times$ $20$ mm $\times$ $20$ mm, serves as the primary domain for measurements. To enable particle rotational motion tracking, a Photron NOVA S6 high-speed camera, operating at 3000 frames per second with a resolution of 1024x1024 pixels, is employed to record particle motion within the flow field.
Magnetic field generator: A uniform rotating magnetic field within a restricted measurement volume, $\bm{H}=H\bm{h}$, with an angular frequency $\omega_{H}$, is generated using a system of two pairs of mutually perpendicular Helmholtz coils. The vertical pair (marked in red in Fig. 1a) has an interior diameter of 300 mm, while the horizontal pair (marked in blue in Fig. 1a) has an interior diameter of 245 mm. For each pair, the separation between the two coils is equal to the corresponding coil diameter.
@@
Magnetic particles: The magnetic particles are produced by coating spherical Styrofoam cores with a layer of magnetic paint, which is paramagnetic. The Styrofoam cores are initially arranged in a fixed, evenly spaced configuration on a substrate. The magnetic paint is then applied uniformly from the top of the substrate to the Styrofoam spheres through a spraying process. Finally, the coated particles are detached from the substrate.
@@
These particles exhibit anisotropy in three key aspects: (1) Shape: The particles are approximately spherical, with a mean diameter of $0.762\pm 0.066$ mm. (2) Density: The magnetic particles have a mean density of $\rho_{\text{particle}}=0.208\pm 0.014$ kg/m3. (3) Magnetic Properties: Due to the non-uniform distribution of the magnetic paint on the particle surface, the magnetic properties vary between particles. Specifically, the magnetic moment magnitude $|\bm{m}|$, the orientation of the magnetic moment, and the spatial distribution of magnetization across the surface differ from particle to particle. This variability influences how individual particles respond to an external magnetic field. All these three aspects of particle polydispersity can introduce minor differences between the simulation results and the experimental observations. However, these differences remain small, and overall, the simulations show good agreement with the experiments. More importantly, our theoretical modeling and the simulations successfully capture key physical features observed in the experiments, such as the location of peak responses.
Particle rotation tracking: The particle rotational motion is tracked using an efficient particle rotation tracking algorithm. The algorithm initiates by pre-processing the experimental images through the application of Gaussian filtering to mitigate noise, contrast enhancement to optimize image quality, and central cropping to ensure precise target particle localization. Subsequently, a set of discrete candidate rotation angles is defined, and corresponding theoretical rotated images are generated utilizing 3D rotation matrices. For each temporal frame, the algorithm computes the cross-correlation coefficient between the actual experimental image and each theoretical rotated image to determine the optimal matching rotation angle. By tracking these optimal angles across the time series, the rotational motion trajectory of the particle is reconstructed and visualized.
@@
A paramagnetic particle moves in a homogeneous and isotropic turbulent flow while subjected to an external magnetic field $\bm{H}$. The magnetic field rotates in the $xoy$-plane (laboratory frame) with an angular velocity $\omega_{H}$ and intensity ${H}$, expressed as $\bm{H}=H\bm{h}=H(\cos(\omega_{H}t),\sin{(\omega_{H}t)},0)^{T}$. The particle is modeled as a sphere with intrinsic magnetic anisotropy (same as that of a prolate spheroid of aspect ratio 1.2), characterized by a preferred magnetization direction $\bm{n}$, which dynamically evolves over time. This instantaneous unit vector is given by $\bm{n}=(\sin\alpha\cos\psi,\sin\alpha\sin\psi,\cos\alpha)^{T}$, where $\alpha$ and $\psi$ (defined in Fig. 2a) describe the orientation of $\bm{n}$.
@@
where $\omega_{a}=\mu\mathcal{V}_{p}H^{2}(\chi_{\parallel}-\chi_{\bot})/\xi_{r}$ is the normalized magnetic field intensity which is also called anisotropic frequency [8, 17], with $\omega_{a}>0$ ($\omega_{a}<0$) for a prolate (oblate) particle. The perpendicular rotational component, $\bm{\omega}_{p,\bot}$, can be obtained by $\bm{\omega}_{p,\bot}=\bm{n}\times\dot{\bm{n}}$ and represents the particle tumbling rate. $\bm{\omega}_{f}$ is half of the vorticity of the background flow field, i.e., $\bm{\omega}_{f}=\frac{1}{2}[\bm{\nabla}\times\bm{u}_{f}]_{O}$ with the subscript O implying evaluation at the center of mass of the sphere, which is obtained through performing Direct Numerical Simulations (DNS) using a pseudo-spectral method and details can be found later.
@@
where for the overdamped particle, we have $\frac{d\left(\xi_{r}\bm{\omega}_{p,\parallel}\right)}{dt}=\bm{0}$.
@@
In a quiescent fluid, the critical frequency at which the particle transitions from synchronous rotation to oscillatory motion is given by: $\omega_{cr}=\frac{\omega_{a}}{2}$. If $\omega_{H}\leq\omega_{cr}$, the particle follows the rotating magnetic field synchronously with a constant phase lag $\beta$. Beyond this threshold, the particle exhibits back-and-forth oscillations [18].
@@
To establish the evolution equation for $\alpha$, we introduce a unit vector $\bm{N}=(-\sin\psi,\cos\psi,0)^{T}$, which denotes the direction perpendicular to $\bm{n}_{xoy}$, the projection of $\bm{n}$ onto the $xoy$-plane (see Fig. 2a for the definition). Projecting Eq. (5) along $\bm{N}$ leads to
@@
Eq. (12) is used to predict the phase transition between the phase-locked regime and the turbulent-dynamics and back-and-forth regimes (blue solid line in Fig. 3m), and is ultimately describing the dynamics that gives rise to the mechanism of stochastic resonance (Figs. 4c-g). Similar dynamical equations as Eq. (12) have been widely applied in multiple physical systems, including condensed matter physics, chemical and biological systems, optics and laser physics, electrical engineering, and financial econometrics [54, 55], with a range of related physical problems, such as Josephson junctions [56, 57, 58, 59, 60], vortex diffusion and flux-line dynamics in superconductors [61, 56, 62], reaction-rate characteristics relevant to chemistry, engineering, and biology [63], Brownian motors [64], and decision-making processes under uncertainty which resemble financial market fluctuations [55, 60]. The universality of this equation suggests that our findings have broad implications beyond the specific context of magnetic particles in turbulence, offering deeper theoretical insights into nonlinear dynamical systems and facilitating the development of new applications in diverse scientific and technological domains.
Now we show the details of predicting the boundary (Fig. 3m, blue solid line) that separates the phase-locked regime (Fig. 3m, purple region) from the back-and-forth and turbulent-dynamics regimes. The phase-locked regime exists if the equation $\dot{\beta}=0$ admits fixed points. Based on the simplified governing equation, Eq. (2) or Eq. (12), this condition translates to ensuring that the following equation has real solutions for any imposed $\omega_{H}$:
@@
The left-hand side of Eq. (13) is bounded within $[-\omega_{a}/2,\omega_{a}/2]$, while the right-hand side varies typically over $\left[\omega_{H}-\sqrt{\sigma(\omega_{f,z})},\omega_{H}+\sqrt{\sigma(\omega_{f,z})}\right]$, where $\sigma(\omega_{f,z})$ is the variance of $\omega_{f,z}$. For Eq. (13) to have real solutions for any imposed $\omega_{H}$, the following condition must hold:
@@
where $[\bm{u}_{f}]_{O}$ denotes the fluid velocity with the subscript $O$ indicating evaluation at the center of the particle, $\bm{x}_{p}$ represents the particle position, and $\bm{v}_{p}=\dot{\bm{x}}_{p}$ is the particle velocity. The Lagrangian vorticity is computed as $\bm{\omega}_{f}=\frac{1}{2}[\bm{\nabla}\times\bm{u}_{f}]_{O}$. The density contrast is defined as $\beta_{p}=\frac{3}{1+2\rho_{p}/\rho_{f}}$, where $\rho_{p}/\rho_{f}$ is the particle-to-fluid density ratio. The Stokes number is given by $\text{St}=d_{p}^{2}/(12\beta_{p}\nu\tau_{\eta})$, where $d_{p}$ is the particle diameter and $\nu$ is the fluid kinematic viscosity.
@@
A key consideration in extending our results to higher Reynolds number turbulence is the effect of intermittency, particularly the heavy-tailed distribution of vorticity fluctuations in fully developed turbulence. In principle, for very large Reynolds numbers, the intermittency of the small-scale vorticity field should be accounted for, as extreme vorticity events become more pronounced. However, previous studies have shown that while intermittency affects the fine-scale structure of turbulence, its influence on the rotational dynamics of small particles (compared to the Kolmogorov length scale), particularly in the parameter regimes relevant to our study, remains relatively weak, because the intense vorticity events, although frequent, are spatially localized and short-lived; thus, their cumulative effect on particle rotation is averaged out [69, 72, 73]. Our results and the underlying mechanisms in the main text provide valuable insights into the fundamental principles governing particle rotation in turbulent flows. Future research could explore the effects of finite-sized particles and beyond the dilute regime.

### v2

@@ Abstract
In strongly subcritical regime ($\omega_{H}/\omega_{a}\ll 1/2$), the resonance peak is purely set by the turbulent noise, making it sensitive to the preferential sampling effect [7, 10, 11] (see Supplementary Information for neutral and heavy particles results). Here we introduce the relaxation time $\tau^{*}=1/\left(\omega_{a}\sqrt{1-4\big{(}\frac{\omega_{H}}{\omega_{a}}\big{)}^{2}}\right)$. Physically, $\tau^{*}$ represents the characteristic time it takes for a perturbed particle rotational motion to return to its phase-locked state, i.e., $\dot{\beta}=0$, after a small disturbance (detailed derivations are in Supplementary Information). Close to the critical ratio $\omega_{H}/\omega_{a}=1/2$, the significantly increasing particle relaxation time $\tau^{*}$ shifts the resonance position, making it the dominant factor. Therefore, a longer $\tau^{*}$ allows particles to escape the potential well even at lower noise levels.
@@
A striking consequence of SR is the symmetry-breaking effect, manifesting as a “zero plus zero is greater than zero” phenomenon. This leverages the fact that in the weak noise regime ($\omega_{\eta}<\omega_{a}$), the response of the system to the noise is non-linear (Fig. 4b, purple-shaded part).
To reveal this effect, instead of applying a constant-frequency rotating magnetic field ($\omega_{H}$), we now impose a time-dependent field rotation with a zero-average, $\omega_{H}(t)$ (Fig. 5a). Each oscillation cycle consists of a forward and a reverse rotation phase, characterized by angular velocities $\omega_{H}^{+}$ and $\omega_{H}^{-}$ with corresponding durations $T_{H}^{+}$ and $T_{H}^{-}$. These parameters are scaled as $\omega_{H}^{+}/\omega_{H}^{-}=T_{H}^{-}/T_{H}^{+}=\gamma$, ensuring that the net angular velocity over a full cycle remains zero. Two distinct response regimes emerge as $\omega_{\eta}/\omega_{a}$ increases for a fixed $\gamma$ (Fig. 5b), highlighting the ability of this oscillation framework to probe the nonlinearity of the system in response to noise. Specifically, the response $\langle\dot{\beta}\rangle/\omega_{\eta}$ decreases monotonically in the nonlinear regime (weak noise limit), while it remains nearly constant and close to zero in the linear regime (strong noise limit, characterized by response collapse, see also Fig. 4b, inset). This behavior aligns well with a superposition model which predicts the response as
@@
Conclusions and outlook
We have revealed the first experimental and numerical evidence of stochastic resonance (SR) for magnetic particles suspended in turbulence, demonstrating how the interplay between turbulent fluctuations and external forcing induces coherent rotational motion of magnetic particles. Three distinct regimes: “phase-locked”, “back-and-forth”, and “turbulent-dynamics” are identified, with SR occurring at the transition between “phase-locked” and “turbulent-dynamics” regimes. Remarkably, even when both turbulence and the magnetic field have zero mean angular velocity, their collective effect is capable of generating a nonzero particle rotational response, as the turbulent vorticity helps the particle to “surf” the rotating magnetic field.
Different particle properties influence their response, and when combined with the stochastic resonance mechanism, they enable a novel magnetic resonance-based approach to measure turbulent vorticity. This method leverages activated magnetic particles as probes, effectively functioning as a turbulent vorticity “microscope” that operates in various conditions. In the present work, we validate this concept using optical measurements of particle rotation. However, in principle, once the angular response is well-characterized, the same protocol could be applied even in optically inaccessible environments by measuring the magnetic fields emitted by the spinning magnetic particle directly. In such scenarios, the activated magnetic particles would function as remotely readable vorticity probes. The protocol is straightforward: by dispersing magnetic particles into turbulence and varying the magnetic frequency $\omega_{H}$ and intensity $\omega_{a}$ while maintaining a constant ratio $\omega_{H}/\omega_{a}$, one can measure the resonance curve of the averaged particle angular response. The location of the resonance peak directly reveals the turbulent vorticity magnitude.
While the present work focuses on understanding particle angular dynamics, it lays essential groundwork for subsequent investigations into how these actively tunable particles can shape turbulent structures,
opening possibilities for flexible manipulation of turbulence, where propeller-like particles could be tuned to navigate specific flow structures, inject energy into turbulence, and influence flow properties such as intermittency [54]. More broadly, our study provides fundamental mechanisms that externally tunable rotational dynamics could be leveraged to design active materials [55, 56, 57]. If organized coherently, their collective behavior may also exhibit hydrodynamic properties analogous to systems with odd viscosity, which may be promising experimental realizations of chiral fluids [58, 59].
@@
Data availability
The data generated during the course of this study is publicly available on Zenodo at https://zenodo.org/records/17076195 [60]
Code availability
The code used for processing the data and generating the figures is publicly available on Zenodo at https://zenodo.org/records/17076195 [60] under the CC-BY-4.0 license.
@@ Methods
The experimental setup comprises three fundamental components: a turbulence generator, a magnetic field generator, and magnetic particles. Detailed specifications of the experimental apparatus, along with a comprehensive description of the particle rotation tracking methodology, are provided in our dedicated instrumentation and methodology paper [61]. Here, we present only the essential information pertinent to the current study.
@@
The working fluid is water, contained within an octagonal-shaped vessel with an internal diameter of $2R=150$ mm and a height of 220 mm. This specific geometry is chosen to enhance optical accessibility for visualization. The flow is driven by two counter-rotating disks, each with a diameter of 150 mm, separated by a distance of 150 mm, and rotating at the same angular velocity magnitude in opposite directions. Each disk is equipped with six straight blades, each 10 mm in height, designed to enhance inertial stirring. The disks are powered by two calibrated DC motors operating at a constant voltage, ensuring that the angular velocity $\Omega$ remains stable over time with a precision of approximately 1%. The motor rotation frequency can be adjusted within the range of $0.25$ to $8$ Hz. Under these conditions, a fully developed turbulent flow with up to $Re_{\lambda}=447$ can be generated in the central region of the container with small-scale statistics close to local isotropy [47, 48, 62, 49, 63, 64, 50]. This region, measuring approximately $20$ mm $\times$ 2020 mm $\times$ 2020 mm, serves as the primary domain for measurements. To enable particle rotational motion tracking, a Photron NOVA S6 high-speed camera, operating at 3000 frames per second with a resolution of $1024\times 1024$ pixels, is employed to record particle motion within the flow field.
Magnetic field generator: A uniform rotating planar magnetic field within a restricted measurement volume, $\bm{H}=H\bm{h}$, with an angular frequency $\omega_{H}$, is generated using a system of two pairs of mutually perpendicular Helmholtz coils. The vertical pair (marked in red in Fig. 1a) has an interior diameter of 300 mm, while the horizontal pair (marked in blue in Fig. 1a) has an interior diameter of 245 mm. For each pair, the separation between the two coils is equal to the corresponding coil diameter.
@@
Magnetic particles: The magnetic particles are produced by coating spherical Styrofoam cores with a layer of magnetic paint, which is paramagnetic. The Styrofoam cores are initially arranged in a fixed, evenly spaced configuration on a substrate. The magnetic paint is then applied uniformly from the top of the substrate to the Styrofoam spheres through a spraying process, ensuring that all particles are coated in a consistent manner; however, at the single-particle level the coating is not perfectly homogeneous, and small variations in thickness or distribution lead to slight differences in magnetic properties between particles. Finally, the coated particles are detached from the substrate.
@@
These particles exhibit anisotropy in three key aspects: (1) Shape: The particles are approximately spherical, with a mean diameter of $0.762\pm 0.066$ mm. (2) Density: The magnetic particles have a mean density of $\rho_{\text{particle}}=208\pm 14$ kg/m3. (3) Magnetic Properties: Due to the non-uniform distribution of the magnetic paint on the particle surface, the magnetic properties vary between particles. Specifically, the magnetic moment magnitude $|\bm{m}|$, the orientation of the magnetic moment, and the spatial distribution of magnetization across the surface differ from particle to particle. This variability influences how individual particles respond to an external magnetic field. All these three aspects of particle polydispersity can introduce minor differences between the simulation results and the experimental observations. However, these differences remain small, and overall, the simulations show good agreement with the experiments. More importantly, our theoretical modeling and the simulations successfully capture key physical features observed in the experiments, such as the location of peak responses.
Particle rotation tracking: The particle rotational motion is tracked using an efficient particle rotation tracking algorithm. The algorithm initiates by pre-processing the experimental images through the application of Gaussian filtering to mitigate noise, contrast enhancement to optimize image quality, and central cropping to ensure precise target particle localization. Subsequently, a set of discrete candidate rotation angles is defined, and corresponding theoretical rotated images are generated utilizing 3D rotation matrices. For each temporal frame, the algorithm computes the cross-correlation coefficient between the actual experimental image and each theoretical rotated image to determine the optimal matching rotation angle. By tracking these optimal angles across the time series, the rotational motion trajectory of the particle is reconstructed and visualized. More details of the tracking methodology are reported in our dedicated instrumentation and methodology paper [61].
@@
A paramagnetic particle moves in a homogeneous and isotropic turbulent flow while subjected to an external magnetic field $\bm{H}$. The magnetic field rotates in the xoyxoy-plane (laboratory frame) with an angular velocity $\omega_{H}$ and intensity H{H}, expressed as $\bm{H}=H\bm{h}=H(\cos(\omega_{H}t),\sin{(\omega_{H}t)},0)^{T}$. The particle is modeled as a sphere with intrinsic magnetic anisotropy (same as that of a prolate spheroid of aspect ratio 1.2), characterized by a preferred magnetization direction $\bm{n}$, which dynamically evolves over time. This instantaneous unit vector is given by $\bm{n}=(\sin\alpha\cos\psi,\sin\alpha\sin\psi,\cos\alpha)^{T}$, where $\alpha$ and $\psi$ (defined in Fig. 2a) describe the orientation of $\bm{n}$.
@@
where $\omega_{a}=\mu\mathcal{V}_{p}H^{2}(\chi_{\parallel}-\chi_{\bot})/\xi_{r}$ is the normalized magnetic field intensity which is also called anisotropic frequency [8, 20], with $\omega_{a}>0$ ($\omega_{a}<0$) for a prolate (oblate) particle.
@@
where for the overdamped particle, we have $\frac{d\left(\xi_{r}\bm{\omega}_{p,\parallel}\right)}{dt}=\bm{0}$. $\bm{\omega}_{f}$ is half of the vorticity of the background flow field, which is obtained through performing Direct Numerical Simulations (DNS) using a pseudo-spectral method and details can be found later. The validity of the overdamped model is further discussed in the Supplementary Information.
@@
To establish the evolution equation for $\alpha$, we introduce a unit vector $\bm{N}=(-\sin\psi,\cos\psi,0)^{T}$, which denotes the direction perpendicular to $\bm{n}_{xoy}$, the projection of $\bm{n}$ onto the xoyxoy-plane (see Fig. 2a for the definition). Projecting Eq. (5) along $\bm{N}$ leads to
@@
Eq. (S.1) is used to predict the phase transition between the phase-locked regime and the turbulent-dynamics and back-and-forth regimes (blue solid line in Fig. 3m), and is ultimately describing the dynamics that gives rise to the mechanism of stochastic resonance (Figs. 4c-g). Similar dynamical equations as Eq. (S.1) have been widely applied in multiple physical systems, including condensed matter physics, chemical and biological systems, optics and laser physics, electrical engineering, and financial econometrics [66, 67], with a range of related physical problems, such as Josephson junctions [68, 69, 70, 71, 72], vortex diffusion and flux-line dynamics in superconductors [73, 68, 74], reaction-rate characteristics relevant to chemistry, engineering, and biology [75], Brownian motors [76], and decision-making processes under uncertainty which resemble financial market fluctuations [67, 72]. The universality of this equation suggests that our findings have broad implications beyond the specific context of magnetic particles in turbulence, offering deeper theoretical insights into nonlinear dynamical systems and facilitating the development of new applications in diverse scientific and technological domains.
Now we show the details of predicting the boundary (Fig. 3m, blue solid line) that separates the phase-locked regime (Fig. 3m, purple region) from the back-and-forth and turbulent-dynamics regimes. The phase-locked regime exists if the equation $\dot{\beta}=0$ admits fixed points. Based on the simplified governing equation, Eq. (2) or Eq. (S.1), this condition translates to ensuring that the following equation has real solutions for any imposed $\omega_{H}$:
@@
The left-hand side of Eq. (13) is bounded within $\omega_{a}/2,\omega_{a}/2]$, while the right-hand side varies typically over $\left[\omega_{H}-\sqrt{\sigma(\omega_{f,z})},\omega_{H}+\sqrt{\sigma(\omega_{f,z})}\right]$, where $\sigma(\omega_{f,z})$ is the variance of $\omega_{f,z}$. For Eq. (13) to have real solutions for any imposed $\omega_{H}$, the following condition must hold:
@@
where $\bm{u}_{f}]_{O}$ denotes the fluid velocity with the subscript $O$ indicating evaluation at the center of the particle, $\bm{x}_{p}$ represents the particle position, and $\bm{v}_{p}=\dot{\bm{x}}_{p}$ is the particle velocity. The Lagrangian vorticity is computed as $\bm{\omega}_{f}=\frac{1}{2}[\bm{\nabla}\times\bm{u}_{f}]_{O}$. The density contrast is defined as $\beta_{p}=\frac{3}{1+2\rho_{p}/\rho_{f}}$, where $\rho_{p}/\rho_{f}$ is the particle-to-fluid density ratio. The Stokes number is given by $\text{St}=d_{p}^{2}/(12\beta_{p}\nu\tau_{\eta})$, where $d_{p}$ is the particle diameter and $\nu$ is the fluid kinematic viscosity.
@@
A key consideration in extending our results to higher Reynolds number turbulence is the effect of intermittency, particularly the heavy-tailed distribution of vorticity fluctuations in fully developed turbulence. In principle, for very large Reynolds numbers, the intermittency of the small-scale vorticity field should be accounted for, as extreme vorticity events become more pronounced. However, previous studies have shown that while intermittency affects the fine-scale structure of turbulence, its influence on the rotational dynamics of small particles (compared to the Kolmogorov length scale), particularly in the parameter regimes relevant to our study, remains relatively weak, because the intense vorticity events, although frequent, are spatially localized and short-lived; thus, their cumulative effect on particle rotation is averaged out [81, 84, 85]. Our results and the underlying mechanisms in the main paper provide valuable insights into the fundamental principles governing particle rotation in turbulent flows. Future research could explore the effects of finite-sized particles and beyond the dilute regime.